Ok this problem is really confusing, I feel like I might know what to do but I cannot remember the formula to do it.
In 1992, the life expectancy of males in a certain country was 61.7 years. In 1997, it was 64.5 years Let E express the life expectancy in year t and let t represent the number of years since 1992.
The linear function E(t) that fits the data is
E(t)=__t+__
Use the function to predict the life expectancy of males in 2005.
E(13)=?

Answer by stanbon(57979)

In 1992, the life expectancy of males in a certain country was 61.7 years. In 1997, it was 64.5 years Let E express the life expectancy in year t and let t represent the number of years since 1992.
---------
You have two points of the form (t,E)
One is (0,61.7) ; this is the intercept for the equation.
-------------
The other is (5,64.5)
-------------------------
slope = (64.5 - 61.7)/(5-0) = 2.8/5 = 0.56
------
The linear function E(t) that fits the data is
E(t)=_0.56t+61.7
------------------
Use the function to predict the life expectancy of males in 2005.
E(13) = 0.56*13 + 61.7 = 68.98 years
========================================
Cheers,
Stan H.